OpenStudy (anonymous):

A square park has a diagonal walkway from one corner to another. If the walkway is about 38 yards long, what is the approximate length of each side of the park?

OpenStudy (anonymous):

\[19*\sqrt{2}\]

OpenStudy (radar):

Let x = length of side of square park. The diagonal would be the \[\sqrt{x ^{2}+x ^{2}}=\sqrt{2x ^{2}}\]

OpenStudy (anonymous):

my answer is correct anyway

OpenStudy (radar):

\[\sqrt{2x ^{2}}=38\]Square both sides of the equation getting: \[2x ^{2}=38^{2}=1444\]\[x ^{2}=722\]\[x=\sqrt{722}\]

OpenStudy (anonymous):

can i sove it a(squared) + b(squared)= c(squared) ??

OpenStudy (radar):

Yes and since it is square, the sides are all equal, then you do not need to have but one unknown.

OpenStudy (radar):

As you can determine from my first post, the two sides are squared, and summed and the square root taken. You must take the square root in order to obtain c otherwise you would have c^2

OpenStudy (anonymous):

My answer was correct and much simpler

OpenStudy (radar):

\[\sqrt{722}=\sqrt{361}\sqrt{2}=19\sqrt{2}\]

OpenStudy (radar):

Simplifying my results, I was attempting to show the steps that were taken to obtain the simple answer.

OpenStudy (radar):

A cryptic answer does not necessarily provide the kind of help a serious student may require.

OpenStudy (radar):

summerflower did you follow with understanding?

OpenStudy (anonymous):

not really... im confused on how you got 772 and square root all that. sorry im just stuck on the pythagorean theorem stuff

OpenStudy (radar):

OK I will draw (show and tell)